How do you factor the binomial 21x^5-14x^4
Answer by lwsshak3(11628)

How do you factor the binomial 21x^5-14x^4
**
f(x)=21x^5-14x^4=0
7x^4(3x-2)=0
7x^4=0
x=0 (multiplicity of 4)
or
3x-2=0
x=2/3